# Holstein mechanism in single-site model with unitary evolution

### v3

@@ III.1 Polaronic shift
III.2 Effect of transformation to the boson operators
Since both the fermion operator and boson operator share the same unitary dynamics, they are not mutually commute unless remove the Hermitian part of the boson operators. The difference of position operator before and after transformation, $\sqrt{2\omega}\Delta x(t):=e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}-(b^{{\dagger}}(t)+b(t))$ which is a diagonalizable defective matrix, is not soly determined by the fermion operator due to the contribution from $[b^{{\dagger}}_{nH}(t)-b_{nH}(t),b^{{\dagger}}_{H}(t)+b_{H}(t)]$. The lowered rank signifies the existence of states insensitive to the perturbations (reminiscent of the dark states). Further, since $e^{S}(b^{{\dagger}}(t)+b(t))^{2}e^{-S}=((b^{{\dagger}}(t)+b(t))+\sqrt{2\omega}\Delta x(t))^{2}$, we have
${\rm Tr}[2\omega\Delta x(t)^{2}]+2{\rm Tr}[(b^{{\dagger}}(t)+b(t))\sqrt{2\omega}\Delta x(t)]=0,$ (11)
where for $\omega\leq 1$, we obtain
${\rm Tr}[2\omega\Delta x(t)^{2}]\sim\left\{\begin{array}[]{ll}0,\ (small\ t)\\ \omega(1-e^{-t})\ (large\ t)\end{array}\right.$ (12)
which is linear in $\omega$ at long time. The exact result of ${\rm Tr}[2\omega\Delta x(t)^{2}]$ is shown in Fig.3. We also obtain its approximated result in appendix. Fluctuations signifying non-local correlations can be seen in Fig.3(d), where we apply another transformation using $S^{\prime}=c^{{\dagger}}(0)c(t)(b^{{\dagger}}(t)-b(t))$, which is nomore a anti-Hermitian operator as long as $t\neq 0$. Nonzero value of ${\rm Tr}[2\omega\Delta x(t)^{2}]$ also implies the squeezing effect from the perturbation, similar to the result of canonical transformation for squeezing with squeezing operator $e^{r(b^{2}-(b^{{\dagger}})^{2})}$.
@@ IV Dynamics of reduced density matrix
As shown in Fig.4, $\rho_{e;11}$ is quite large and increase with time, thus $\rho_{e}$ and $\rho_{b}$ are close to purity state and the purity increase with time. This is consistent with the increased purity of each subsystem as shown in Fig.5(d) and the continously decrease of internal entanglement (negativity) of $\rho_{tot}$ as shown in Fig.6(b). Thus the nonMarkovian-to-Markovian transition can be well observed at least before the critical time $t_{c}$ of the emergent degenerated ground state of $H$, during which the evolution is unitary and the system is nearly closed such that as the total density getting thermalized by the environment with lowered internal entanglement, the subsystems have increased purity until the entanglement with the environment sets in (the begining of nonunitary evolution). In other word, in early stage, the weak nonMarkovianity leads to power-law monotonous decay; while in late stage, the degenerate ground state of $H$ cause additional strong nonMarkovianity with significant coherent exchange of information between system and bath, which outweights the Markovian thermal process despite the decoherence dominates at the long-time and implying the enhanced entanglement. Note that the Markovian dynamics in the late stage can still be observed in term of the observables regardless of the density operators. Thus, the perturbation, which is initially well-defined and coherent, in the late stage becomes random and is composed of many unobserved degrees of freedom due to the degeneracy of the system’s ground state.
@@
The dissipative phase transition is only observable in the early stages of evolution. During this period, the dissipation-induced gain and loss can be ignored, and the total Hamiltonian maintains unitary evolution. By setting a finite threshold value for the ground state, Fig.5 presents the scenario where the total density matrix is considered to be in a mixed state. In this case, once time exceeds a critical value, the total system’s density matrix becomes mixed. This implies that the system is indeed an open system and its evolution becomes non-unitary at long times. Here, strong dissipation drives the system toward a state akin to a highly fluctuating thermal state in a closed system (see Fig.5(b)). This observation is consistent with the opening of the dissipative gap in the later stage of evolution. At this point, quantum fluctuations (under the zero-temperature condition) make continuous symmetry breaking and the formation of an ordered phase less likely.
@@
Figure 5: (a) The four diagonal elements (populations) of $\rho_{b}$ for unitary evolution and pure total density, obtained by setting an infinitesimally small threshold value for judging the degenerate ground state. (b)-(c) Populations of $\rho_{b}$ and $\rho_{e}$ when a finite threshold value is used. The non-unitary evolution and mixed total density appear at a critical time $t_{c}\approx 800$, as indicated by the vertical cyan dashed line. (d) Purities of $\rho_{tot}$ (blue) and $\rho_{e}$ (or $\rho_{b}$; red) for finite threshold value (left axis) and infinitesimally threshold value (right axis). The green dashed line indicates when the population starts to deviate from its logarithmic increase, i.e., the emergence of Markovian process. The left red arrow indicates the lowered purity due to the entanglement between subsystems, while the right red arrow indicates the lowered purity due to the entanglement between the subsystem and the bath. Obviously, the effect of the latter dominates over the fomer. That is why the nonMarkovian-type fluctuation overweight the Markovian dynamics in the late stage.
@@ V Conclusion
[…]
The unitary dynamics preserves the phase-space distribution during the Markovian process wth slow relaxation, and the electron-phonon interaction dominates over the influences from bath electrons. In this limit, the Gaussian wave function fail to describe bosons which can no longer be treated as classical field, and phononic time scale is much faster than the electronic time scale. The mean-field theory also fails to be applied in this case. The unitary dynamics as well as boson-frequency-dependence provides the evidence of quantum phase transition from the distinct behaviors at short-time and long-time stages, which exhibit power law and exponential law decay, respectively, as can be seen from the polaronic shift, boson energy, and the dynamics of reduced density $\rho_{e}$. Moreover, the effective mass renormalization $m^{*}/m(=e^{\Sigma(t)/\omega})$ should follow the same rule, despite not being shown here. The gradual opening of the dissipative gap can be further verified by estimating the Lindbladian’s spectrum. This estimation, achievable through Lanczos iteration or the quantum trajectory method, as well as the case of boson squeezing ($(b^{{\dagger}})^{2}+b^{2}$) and the deviation away from minimum uncertainty condition (entanglement increase the classical uncertainty), will be the scope of our next work.

### v4

@@ III.1 Polaronic shift
III.2 Effect of transformation on the boson operators
Since the generator of the transformation $S$ involves the boson operators, it does not commute with the position operator unless the coupling vanishes. We define the shift in the position operator induced by the transformation as
$\sqrt{2\omega}\Delta x(t):=e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}-(b^{{\dagger}}(t)+b(t)).$ (14)
This operator difference arises from the non-commutativity between the generator $S$ and the field operators. Since the transformation is unitary, the trace of the squared position operator must be invariant, i.e., ${\rm Tr}[(e^{S}x(t)e^{-S})^{2}]={\rm Tr}[x(t)^{2}]$. Substituting $e^{S}(b^{\dagger}+b)e^{-S}=(b^{\dagger}+b)+\sqrt{2\omega}\Delta x(t)$ into this invariance condition, we derive the following identity
${\rm Tr}[2\omega\Delta x(t)^{2}]+2{\rm Tr}[(b^{{\dagger}}(t)+b(t))\sqrt{2\omega}\Delta x(t)]=0.$ (15)
This relation allows us to quantify the magnitude of the polaronic shift dynamically. For the regime $\omega\leq 1$, the trace of the squared shift exhibits distinct asymptotic behaviors
${\rm Tr}[2\omega\Delta x(t)^{2}]\sim\left\{\begin{array}[]{ll}0,&(t\ll 1)\\ \omega(1-e^{-t}),&(t\gg 1)\end{array}\right$. (16)
which becomes linear in $\omega$ at long times. The exact numerical result of ${\rm Tr}[2\omega\Delta x(t)^{2}]$ is shown in Fig.3. An approximate analytical form is also provided in the Appendix.A.
In Holstein model and driven harmonic oscillators, $\omega=1$ here represents a resonance point where the phonon frequency matches the electronic or driving energy scale. The expectation value of a Hermitian observable like position squared ($\Delta x^{2}$) must be real. The emergence of an imaginary part for $\omega>1$ implies that the polaron ansatz (the transformation $S$) breaks down or becomes unstable in high-frequency regime. This indicates a phase transition. Below $\omega=1$, a stable polaron forms (real displacement). Above $\omega=1$, the system enters a delocalized or radiative regime where the static polaron picture fails and the complex values indicating the decay or instability of that specific bound state solution.
Fluctuations signifying the non-local correlations can be observed in Fig.3(d). Here we consider a modified transformation generator $S^{\prime}=c^{{\dagger}}(0)c(t)(b^{{\dagger}}(t)-b(t))$. Unlike the static $S$, this operator $S^{\prime}$ is no longer strictly anti-Hermitian for $t\neq 0$ due to the time-dependence of the electron operators in the Heisenberg picture. Consequently the nonzero value of ${\rm Tr}[2\omega\Delta x(t)^{2}]$ in this context implies a deformation of the phase space distribution, analogous to the squeezing effect generated by canonical transformations of the form $e^{r(b^{2}-(b^{{\dagger}})^{2})}$, as a time-dependent transformation often induces Rabi-like oscillations or squeezing effects (periodic expansion and contraction of phase space), which looks exactly like the ringing shown here. The rapid oscillations seen in Fig.3(d) are the signature of non-local correlations and squeezing or breathing modes in the boson field. Unlike the smooth saturation of the standard shift (Fig.3(a)), squeezing induces time-dependent variances that oscillate at $2\omega$. Thus the nonzero value of $\text{Tr}[2\omega\Delta x(t)^{2}]$ implies a deformation of the phase space distribution. This is analogous to the squeezing effect, where the vacuum fluctuation circle is distorted into an ellipse. As this deformed state undergoes free evolution (rotation in phase space at frequency $\omega$), its projection onto the position axis (variance) expands and contracts twice per revolution. This generates the characteristic oscillations at $2\omega$ (twice the natural phonon frequency) observed in Fig.3(d).
@@ IV Dynamics of reduced density matrix
As shown in Fig.4, $\rho_{e;11}$ is quite large and increase with time, thus $\rho_{e}$ and $\rho_{b}$ are close to purity state and the purity increase with time. This is consistent with the increased purity of each subsystem as shown in Fig.5(d) and the continously decrease of internal entanglement (negativity) of $\rho_{tot}$ as shown in Fig.6(b). Thus the nonMarkovian-to-Markovian transition can be well observed at least before the critical time $t_{c}$ of the emergent degenerated ground state of $H$, during which the evolution is unitary and the system is nearly closed such that as the total density getting thermalized by the environment with lowered internal entanglement, the subsystems have increased purity until the entanglement with the environment sets in (the begining of nonunitary evolution). In other word, in early stage, the weak nonMarkovianity leads to power-law monotonous decay; while in late stage, the degenerate ground state of $H$ cause additional strong nonMarkovianity with significant coherent exchange of information between system and bath, which outweights the Markovian thermal process despite the decoherence dominates at the long-time and implying the enhanced entanglement. Note that the Markovian dynamics in the late stage can still be observed in term of the observables regardless of the density operators. Thus, the perturbation, which is initially well-defined and coherent, in the late stage becomes random and is composed of many unobserved degrees of freedom due to the degeneracy of the system’s ground state. From Fig.5(d), we also notice that $\rho_{b}$ and $\rho_{tot}$ have richer information exchange, in the late stage, compares to the $\rho_{e}$, as manifested in stronger fluctuation of purity, which implies that, even for a separable multi-qubit system (see Fig.6(b)), the remaining classical correlations can still affect the exchange of information with the environment.
The dissipative phase transition is only observable in the early stages of evolution. During this period, the dissipation-induced gain and loss can be ignored, and the total Hamiltonian maintains unitary evolution. When the ground state is resolved with infinite precision (infinitesimal threshold), the total density matrix $\rho_{tot}$ remains a pure state $|\psi(t)\rangle\langle\psi(t)|$ governed by the Schrödinger equation. In this ideal limit, the system is closed, and entropy is generated solely in the subsystems ($\rho_{e},\rho_{b}$) via entanglement between the electron and the phonon bath.
@@
When we introduce a finite threshold value to define the ground state manifold, a transition occurs at a critical time $t_{c}$ where the ground state becomes highly degenerate. For $t>t_{c}$, the degeneracy introduces an effective disorder. The evolution is no longer well-described by a single trajectory but requires a statistical mixture over the degenerate manifold, $\rho_{tot}=\frac{1}{d}\sum_{j=1}^{d}|\psi_{0;j}\rangle\langle\psi_{0;j}|$. This coarse-graining acts as an information loss mechanism. As noted in Appendix B, this is equivalent to considering the perturbation as inherently random, which induces dephasing and mixing. Thus by setting a finite threshold value for the ground state, Fig.5 presents the scenario where the total density matrix is considered to be in a mixed state. Once time exceeds a critical value, the total system’s density matrix becomes mixed, implying that the system is indeed an open system and its evolution becomes non-unitary at long times. Here, strong dissipation drives the system toward a state akin to a highly fluctuating thermal state in a closed system (see Fig.5(b)). This observation is consistent with the opening of the dissipative gap in the later stage of evolution. At this point, quantum fluctuations (under the zero-temperature condition) make continuous symmetry breaking and the formation of an ordered phase less likely. Consequently, in finite-threshold regime, the total system effectively behaves as an open system. The "environment" here is identified with the unobserved internal degrees of freedom within the degenerate ground state manifold and the stochastic fluctuations of the perturbation. Strong dissipation drives the system toward a state akin to a highly fluctuating thermal state (see Fig.5(b)). This leads to the unitary-to-nonunitary transition observed in Fig.6(b), where the total system entropy increases (purity decreases) after $t_{c}$, due to the entanglement with the environment where the environment is the effective bath created by the disorder and degeneracy. This observation is also consistent with the opening of a dissipative gap in the later stage of evolution, where the quantum fluctuations (under the zero-temperature condition) suppress continuous symmetry breaking, making the formation of an ordered phase less likely.
@@
Figure 5: (a) The four diagonal elements (populations) of $\rho_{b}$ for unitary evolution and pure total density, obtained by setting an infinitesimally small threshold value for judging the degenerate ground state. (b)-(c) Populations of $\rho_{b}$ and $\rho_{e}$ when a finite threshold value is used. The non-unitary evolution and mixed total density appear at a critical time $t_{c}\approx 800$, as indicated by the vertical cyan dashed line. (d) Purities of $\rho_{tot}$ (blue), $\rho_{e}$ (red), and $\rho_{b}$ (green) for finite threshold value (left axis) and infinitesimally threshold value (right axis). The green dashed line indicates when the population starts to deviate from its logarithmic increase, which signifies the emergence of a Markovian process. The left red arrow indicates the lowered purity due to entanglement between subsystems, while the right red arrow indicates the lowered purity due to entanglement between the subsystem and the bath. Obviously, the effect of the latter dominates the former. That is why non-Markovian type fluctuations outweigh the Markovian dynamics in the late stage.
@@ V Conclusion
We investigated electron-phonon coupling in a single-site system characterized by unitary evolution intrinsic to the boson operators within the nonadiabatic limit. We focused on the regime where the phase-space distribution retains its Gaussian nature. Specifically, the variances of position and momentum satisfy the minimum uncertainty relation $\langle x^{2}\rangle=(4\langle p^{2}\rangle)^{-1}$, characteristic of a coherent state (or displaced vacuum) at zero temperature. This minimum uncertainty is guaranteed by the fact that the reduced density matrices remain nearly pure states during the evolution, i.e., ${\rm Tr}[\rho_{e}^{2}]={\rm Tr}[\rho_{b}^{2}]\approx 1$. Since the total evolution is unitary and the global state remains pure (${\rm Tr}\rho_{tot}^{2}=1$), this high purity implies weak entanglement in the bipartite system, although entanglement is slightly enhanced at long times due to the emergence of Markovian dynamics and decoherence processes.
This observed transition from non-Markovian to Markovian dynamics, accompanied by the crossing from a zero dissipative gap to an open dissipative gap, implies a change in the scaling of correlations. This behavior is analogous to the transition from an area law to volume law scaling of entanglement, or the transition from higher to lower spatial dimensions with an enhanced role of dissipation. We also note that this transition can be viewed as continuous in time: shifting from a time-dependent Lindbladian (governing the non-Markovian power-law decay stage) to a time-independent Lindbladian (governing the Markovian exponential decay stage), provided the total density matrix describes a pure state. The evolution of the reduced state, formally written as $|\rho_{e}(t)\rangle\rangle=\mathcal{T}e^{\int^{t}_{0}\mathcal{L}(\tau)d\tau}|\rho_{e}(0)\rangle\rangle$, is analogous to continuous connected gapped Liouvillians in systems of finite spatial dimension mediated by long-range interactions[11].
The unitary dynamics preserves the structure of the phase-space distribution during the Markovian process, characterized by slow relaxation where the electron-phonon interaction dominates over the influence of the phonon bath. In this nonadiabatic limit, the simple Gaussian wave function approximation fails to fully capture the bosons, which can no longer be treated purely as a classical field, and the phononic time scale becomes comparable to or faster than the electronic time scale. Thus standard mean-field theory also fails in this regime.
Our results provide evidence of a dynamical phase transition, evidenced by distinct behaviors at short-time and long-time stages. The unitary dynamics and the boson-frequency dependence lead to a crossover from power-law decay to exponential decay, as observed in the polaronic shift, boson energy, and the dynamics of the reduced density matrix $\rho_{e}$. Moreover, the effective mass renormalization $m^{*}/m(=e^{\Sigma(t)/\omega})$ is expected to follow the same scaling rule. The gradual opening of the dissipative gap can be further verified by estimating the spectrum of the effective Lindbladian. This estimation, achievable through Lanczos iteration or quantum trajectory methods—as well as the investigation of boson squeezing ($(b^{{\dagger}})^{2}+b^{2}$) and deviations from the minimum uncertainty condition driven by entanglement—will be the scope of our future work.